OpenStudy (anonymous):

find solution of differential equation. I really need help, not sure what to do!! 4y''-y=xe^(x/2)

OpenStudy (anonymous):

I have the general solution, i just dont know how to get the particular solution with variation of parameters

OpenStudy (theeric):

Hi! I'm just learning ODE now! But I know you can't get a particular solution without given conditions. You'd need something like \(y(t_0)=y_0\qquad\)and\(\qquad y'(t_0)=y_0'\)

OpenStudy (theeric):

So the general solution would be what you want, I think. It would be a "family of solutions," meaning a bunch of solutions that are related. Get it? Related... Family.. Hehe... Anyway, you would end up with constants. And you'd have to have values for them for the solution to be specific.

OpenStudy (theeric):

Does that sound about right?

OpenStudy (anonymous):

I got the homogeneous solution: \[y=c_{1}e ^{x/2}+c _{2}e ^{-x/2}\] I just need the particular solution to satisfy (xe^(x/2))/4. I do not need initial values for this problem

OpenStudy (theeric):

Ohh... Okay, I haven't learned this yet. This looks like characteristic equation or principle of superposition stuff, but I don't know if I can handle this problem, sorry!

OpenStudy (theeric):

Why don't you take the derivatives of \(y\) and plug them in?

OpenStudy (anonymous):

do you know of anyone who could help?

OpenStudy (theeric):

That didn't help.. Here's my work... It didn't help find a specific solution. \(y=c_{1}e ^{x/2}+c _{2}e ^{-x/2}\) \(y'=\frac{1}{2}c_1e^{x/2}-\frac 1 2 c_2e^{-x/2}\) \(y''=\frac{1}{4}c_1e^{x/2}+\frac 1 4 c_2e^{-x/2}\) Since \(4y''-y=xe^{x/2}\), we substitute in \(y\) and \(y''\) to find that \(4\left(\frac{1}{4}c_1e^{x/2}+\frac 1 4 c_2e^{-x/2}\right)-\left(c_{1}e ^{x/2}+c _{2}e ^{-x/2}\right)=xe^{x/2}\) \(\left(c_1e^{x/2}+c_2e^{-x/2}\right)-\left(c_{1}e ^{x/2}+c _{2}e ^{-x/2}\right)=xe^{x/2}\) \(c_1e^{x/2}+c_2e^{-x/2}-c_{1}e ^{x/2}-c _{2}e ^{-x/2}=xe^{x/2}\) \(0=xe^{x/2}\)

OpenStudy (anonymous):

do you know of anyone who could help??

OpenStudy (theeric):

I messaged somebody on here!

OpenStudy (anonymous):

can you direct them to this problem. I need to turn it in soon. running out of time, AH!

OpenStudy (theeric):

I gotcha!

OpenStudy (theeric):

What process do you solve it by? If \(xe^{x/2}\) was \(0\), I would use the characteristic equation stuff.

OpenStudy (anonymous):

The problem requires Variation of Parameters.

OpenStudy (theeric):

But I don't think the characteristic equation stuff applies if that's not a 0.

OpenStudy (fibonaccichick666):

is y a function of x or t?

OpenStudy (anonymous):

it does, there are just additional steps (wronskian...) to solve for the particular solution that satisfies the nonhomogeneous part on the right hand side.

OpenStudy (theeric):

Haha, I put \(t\)'s there out of habit... Must be \(x\)!

OpenStudy (anonymous):

x

OpenStudy (fibonaccichick666):

k so the other side

OpenStudy (theeric):

Okay, I will be getting to the Wronskian stuff Friday or next week... Not in time for this... Unless I learn very quickly.

OpenStudy (fibonaccichick666):

Wronski isn't bad, it's just a matrix to test LI

OpenStudy (fibonaccichick666):

so sshot, try and test your solution, is it LI? I'm guessing not, because you need an \(x*e^{x/2}\), which you can get from them not being LI and adding in the x. Then you should get a good answer

OpenStudy (fibonaccichick666):

Check example 5 from that link

OpenStudy (anonymous):

Ok, I think i got it now.

OpenStudy (fibonaccichick666):

np, best of luck!

OpenStudy (theeric):

Thanks, @FibonacciChick666 !